OpenStudy (cenaida):

Write the equation of the line shown in the graph in slope-intercept form. line going through points negative 2 comma 0 and 0 comma 5 A. y = (2/5)x – 2 B. y = (2/5)x + 5 C. y = (5/2)x – 2 D. y = (5/2)x + 5

OpenStudy (shamil98):

go to sleep cen its 11 pm

OpenStudy (solomonzelman):

(-2,0) (0,5) the equation is given: Slope: \[\huge\color{red}{\frac{y_1-y_2}{x_1-x_2}}\] \[\huge\color{red}{\frac{0-5}{-2-0}}\]

OpenStudy (cenaida):

naah its still eraly 11 isnt even late haha

OpenStudy (shamil98):

then again bedtime for online schoolers doesn't really matter.

OpenStudy (solomonzelman):

\[\huge\color{blue}{\frac{-5 }{ -2}~~~~->~~~~~\frac {5 }{ 2}}\]SLOPE

OpenStudy (anonymous):

D. y = (5/2)x + 5

OpenStudy (solomonzelman):

Wrong!

OpenStudy (solomonzelman):

y=5/2 x + b since you know it goes through (0,-5) you see (0,-5) is the y-intercept, right? y=5/2 x -5

OpenStudy (anonymous):

Its (0,5)

OpenStudy (solomonzelman):

My bad, that you are right.

OpenStudy (solomonzelman):

@hanastic, really sorry, I made an error...

OpenStudy (solomonzelman):

it's D, as he said.

OpenStudy (cenaida):

Ok thank you hanastic :)